fimd the sum of the first 30 numbers:6+10+14+18+22
Answer by Jstrasner(112)

Ok so first this is an arithmetic equation. To find the sum, you would use this formula:
S(n)=(n/2)(a1+a(n))
a1 being the first number in the sequence.
To find the 30th term, we use the definition of a arithmetic series:
a(n)= a1+(n-1)r => a(30)=6+29(4)=122
Therefore:
s(30)=(30/2)(6+122)= 1920
i hope this helps!